OpenStudy (anonymous):

Joining the frequent visitor program at the planetarium costs $5 per year, Write an equation that can be used to determine n, the number of visits per year for which the cost of being a member of the frequent visitor program is equal to the cost of not being a member

OpenStudy (anonymous):

@nincompoop @adrynicoleb @Kayla_Is_Smart @kayla_renae @kaylaprincess @bass53 Please help, don't get this :(

OpenStudy (anonymous):

@Destinymasha Can you help me please? This is due in a couple hrs :(

OpenStudy (anonymous):

I am not an math expert at this but I will give it a try.

OpenStudy (anonymous):

The cost for a student who is a member of the planetarium's frequent visitor program is $11, a student who is not in the program has to pay $13

OpenStudy (anonymous):

I think the equation is 5y + 11n = 13n Y stands for each year of the membership, n is the number of visits

OpenStudy (anonymous):

I need to find the point in which the cost is the same for members and non members

OpenStudy (anonymous):

@hartnn @Whitemonsterbunny17 @tHe_FiZiCx99 @skullpatrol @inkyvoyd @bibby @midhun.madhu1987 @Conqueror Sorry for mass tagging, I am desperate and I really need this done :( Please help!

OpenStudy (anonymous):

@midhun.madhu1987 can you assist me, for some reason I just can't get this problem down :/

OpenStudy (anonymous):

Ok I found y! y = 0.4 5(0.4) + 11n = 13n 2n + 11n = 13n 13n = 13n But how does this solve my question :/

OpenStudy (midhun.madhu1987):

No idea.. :(

OpenStudy (anonymous):

Do you think something is missing from my text @midhun.madhu1987 as I'm not sure if I wrote down all the information?

OpenStudy (anonymous):

I'm doing a series of questions that had two different main bodies of information. I will list the other one

OpenStudy (anonymous):

A group of 19 students want to see the show at the planetarium. Tickets cost $11 for each student who is a member of the planetarium's frequent visitor program and $13 for each student who is not a member. The total cost of the student's tickets is $209. Use this information for items 9-16.

OpenStudy (anonymous):

I solved that one now I'm onto the second part which is questions 13-16

OpenStudy (anonymous):

What is the part that you don't understand @midhun.madhu1987 I'm not sure what I'm supposed to do

OpenStudy (kaylaprincess):

i'm horrible at math. I do bio sutff ^.^